Attended: Earl, Warren, Prad, Mas, Momtaz
* Web cameras for Hobart 26m not updating. Warren says that he'll ask brett about it. * Warrens says observers please make sure that experiment starts on time by using delay start commands.
* error m5 -900 not while recording or playing for AUA050. Scheduling issue , mostly when two scan have very less time between them. * Large disk offset for Katherine. Lucas says probably because of bad scheduling. again scheduling issue. probably have to move to next field system when Jamie gets back. * dbbc=pps_delay isn't stable at Katherine and changing from 61971 to 61950. Warren says for now, try to reconfigure dbbc between experiments.
Yarragadee 12m: eRemote control is not working since thrusday morning (21 March). for now, use Vncviewer on pcfsyg for observing.